6x-1 over 5, minus, 2x over 3, equals 3. I am solving for x. I tried using the LCM of 15 but every answer I come up with doesn't equal out.
Answer by Juhi(44)

3(6x-1)-5(2x)=3
18x - 3 - 10x = 3
18x - 10x = 3 + 3
8x = 6
x = 6/8
= 3/4